# One protein is all you need

### v2

@@ Appendix D Implementation details
Optimization.
We minimize the loss defined in Equation˜2 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (Ruder, 2016). While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model (Hayes et al., 2024; Lin et al., 2023). Moreover, many models are pre-trained using the Adam optimizer (Kingma & Ba, 2015) or its variants (Loshchilov & Hutter, 2019). However, it was shown that Adam results in less predictable behavior of test-time training compared to the SGD optimizer, possibly due to its more exploratory behavior (Gandelsman et al., 2022).
Customizing large models.
We aim for customization to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead. Since state-of-the-art models for many protein-oriented tasks are typically large, with up to billions of parameters, our aim presents two key challenges. First, when using pre-trained Transformers on a single GPU, even for the forward pass, the batch size is typically limited to only several samples due to the quadratic complexity of the inference (Vaswani, 2017). Second, for the backward pass, even a batch size of one is not always feasible for large models. To address the first challenge, we perform forward and backward passes through a small number of training examples and accumulate gradients to simulate updates with any batch size. We address the second challenge by employing low-rank adaptation (LoRA; Hu et al. (2021)), which in practice enables fine-tuning of any model for which a forward pass on a single sample is feasible, due to a low number of trainable parameters. Section˜G.3 details how ESMFold (Lin et al., 2023), with its 3B-parameter ESM2 backbone $f$, can be efficiently customized, retaining its speed advantage while enhancing performance.
Listing 2: Implementation of ESM2 + ProteinTTT within the proteinttt package.
@@
Appendix E Experimental details
In this section, we describe the proposed benchmark suite for the three customization tasks considered in this work: protein structure prediction (Section˜E.1), protein fitness prediction (Section˜E.2), and protein function prediction (Section˜E.3). Each subsection describes the application of ProteinTTT to the respective models, along with details on the data, metrics, and models. Table˜A3 additionally summarizes the hyperparameters used for the application of ProteinTTT to individual models.
E.1 Protein structure prediction
E.1.1 Datasets
@@
To evaluate the capabilities of ProteinTTT on protein structure prediction, we employ the CAMEO validation and test sets as described in Lin et al. (2023). Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^4] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022. Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023). Therefore, we subselect the challenging validation and test sets where customization with ProteinTTT is suitable.
Specifically, we apply two standard criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors. Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al. (2023)). After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively.
E.1.2 Metrics
To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013).
@@ E.1.3 Models
Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention. This enables fine-tuning ESMFold + ProteinTTT on a single GPU.
@@
Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism for sampling multiple structure predictions. Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al. (2023)). To sample more predictions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids before passing them to the language model. Selecting the solution with the highest pLDDT may lead to improved predicted structure. Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + ProteinTTT, we employ the former as a baseline, running the method for the same number of steps.
@@
Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (Devlin, 2018), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens). The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (Razavi et al., 2019) operating on the protein geometry. In our experiments, we use the smallest, publicly available version of the ESM3 model (ESM3_sm_open_v0) [^6] https://github.com/evolutionaryscale/esm.
@@
HelixFold-Single.
HelixFold-Single is an MSA-free protein structure prediction model that combines representations from a pretrained protein language model with adapted AlphaFold2 geometric modules (EvoformerS and Structure) to directly predict atomic coordinates (Fang et al., 2023). We use the official implementation [^7] https://github.com/PaddlePaddle/PaddleHelix/tree/dev/apps/protein_folding/helixfold-single
HelixFold-Single + ProteinTTT.
E.2 Protein fitness prediction
E.2.1 Datasets
@@
ProteinGym [^8] https://github.com/OATML-Markslab/ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024). The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins. We focus on the well-established zero-shot setup of the benchmark and do not experiment with the supervised setup, as it has not yet been fully incorporated into the official codebase at the time of this study. In total, the dataset contains 2.5M mutants with annotated ground-truth fitness. Since ProteinGym does not contain a data split for the zero-shot setup, employed in this work, we use the whole dataset as the test set.
@@
Figure A3: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work. A) MaveDB, mined from Esposito et al. (2019), includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset. This is largely because most MaveDB assays post-filtering date to 2024, whereas the latest assays in ProteinGym date to 2023. B, C, D) MaveDB is of sufficient quality for model evaluation. Representative baselines, ESM2 and SaProt with both 35 million and 650 million parameters, evaluated on ProteinGym generalize effectively to MaveDB, following a similar distribution of predictions. Panel D illustrates the random subset of 50 proteins used for hyperparameter tuning for fitness prediction. Each point in the plots represents one protein and shows the Spearman correlation averaged across all assays corresponding to the protein (typically one assay per protein). The box plots standardly depict quartiles, medians, and outliers.
To establish a validation set disjoint from ProteinGym (Notin et al., 2024), we mined MaveDB [^9] https://www.mavedb.org (Esposito et al., 2019). As of August 1, 2024, the database contains 1178 Multiplexed Assays of Variant Effects (MAVEs), where each assay corresponds to a single protein, measuring the experimental fitness of its variants. We applied quality control filters to remove potentially noisy data. Specifically, we ensured that the UniProt identifier (Consortium, 2023) is valid and has a predicted structure available in the AlphaFold DB (Varadi et al., 2022). We also excluded assays with fewer than 100 variants, as well as those where at least one mutation had a wrongly annotated wild type or where most mutations failed during parsing. Additionally, to ensure no overlap between datasets, we removed any assays whose UniProt identifier matched with those in ProteinGym, ensuring that the validation and test sets contain different proteins.
@@ E.2.2 Metrics
When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (Notin et al., 2024). First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels. Then, we average the scores across five categories of assayed phenotypes, measuring the effects of mutations: catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
@@
Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table˜A4). Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from Hopf et al. (2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1<N_{eff}/L<100$, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
E.2.3 Models
@@
The ESM2 model is a bidirectional, BERT-like (Devlin, 2018) Transformer trained on millions of protein sequences using masked modeling (Lin et al., 2023). The goal of protein fitness prediction is to predict the effects of mutations, and PLMs are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021). Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
@@
where the sum iterates over mutated positions $i\in T$ with $p(x^{m}_{i}|x_{\setminus i})$ and $p(x_{i}|x_{\setminus i})$ denoting the predicted probabilities of the mutated amino acid and the original one (i.e., wild type), respectively. The conditionals $x_{\setminus i}$ indicate that the input sequence to the model has the position $i$ masked. In this setup, the native (unmutated) sequence, where $T=\emptyset$, has a predicted fitness of 0. Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations. We follow the ProteinGym benchmark and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2. We use the implementation of ESM2 from ProteinGym.
@@
We also experiment with a structure-aware protein language model, SaProt (Su et al., 2023). SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures. Specifically, SaProt uses the same Transformer architecture but expands its vocabulary by combining the 20 standard amino acid tokens with 20 structural tokens from the 3Di vocabulary, increasing the total alphabet size to 400. The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (Razavi et al., 2019), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van Kempen et al., 2022).
Since SaProt is also a protein language model, it also uses Equation˜5 to score variants. However, please note that SaProt, as implemented in ProteinGym (Notin et al., 2024), uses a slightly different version of the log odds ratio. In SaProt, the conditions in the log probabilities in Equation˜5 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions. During customization with ProteinTTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup. We use the implementation of SaProt from ProteinGym8.
@@
Finally, we experiment with the state-of-the-art fitness predictor, ProSST (Li et al., 2024). ProSST primarily improves upon SaProt (Su et al., 2023) by incorporating a larger vocabulary of structural tokens and employing disentangled attention mechanisms. Instead of relying on the 3Di alphabet optimized for protein structure search with Foldseek (van Kempen et al., 2022), Li et al. (2024) pre-train a new autoencoder to denoise corrupted protein backbones and cluster the resulting latent space using the $K$-means algorithm (Lloyd, 1982). Notably, optimal performance for fitness prediction is achieved with $K=2048$ tokens, compared to just 20 in the 3Di vocabulary used by SaProt. We adopt this model in our experiments. Additionally, disentangled attention in ProSST enhances information propagation between sequence and structure within its Transformer blocks, further improving prediction performance. The model has 110M parameters in total.
ProSST, similarly to ESM2 and SaProt, is pre-trained using masked language modeling applied to protein sequence tokens. To score mutations on the ProteinGym benchmark (Notin et al., 2024), ProSST also uses the log-odds ratio, but in a slightly different way compared to ESM2 and SaProt. Specifically, ProSST performs a single forward pass to predict log probabilities, which are then used to score all mutations. Formally, this approach modifies the log probability condition in Equation˜5, replacing $x_{\setminus i}$ with $x$.
@@
Finally, we experiment with MSA Transformer for fitness prediction (Rao et al., 2021). Similar to ESM2 (Lin et al., 2023), MSA Transformer is pre-trained on large protein sequence datasets; however, it is trained on multiple sequence alignments (MSAs) rather than individual sequences.
Since MSA Transformer is also a protein language model, it can be used for fitness prediction in the same way as ESM2, as discussed above, by computing the log-odds ratio over the first sequence in the MSA in this case. We reproduce the results of MSA Transformer on the ProteinGym benchmark with two modifications: (1) we sample a weighted subset of 32 sequences from each MSA instead of 400, and (2) we use only one random seed instead of five for ensembling. These changes significantly reduce computational time while also slightly improving performance compared to the results reported in ProteinGym. This improvement may be explained by the fact that the performance of MSA Transformer saturates with increasing MSA input size (Figure 4 in Rao et al. (2021)).
@@
E.3 Protein function prediction
E.3.1 Datasets

### v3

@@ Appendix E Implementation details
Code snippet 2: Implementation of ESM2 + ProteinTTT within the proteinttt package.
@@
Infrastructure.
Source code.
We provide a user-friendly and easily extensible PyTorch (paszke2019pytorch) implementation of ProteinTTT, available as the proteinttt Python package  [^3] https://github.com/anton-bushuiev/ProteinTTT . We provide Code snippet˜1 and Code snippet˜2 in Python to demonstrate the implementation of inference and customization with ProteinTTT, respectively. Code snippet˜1 demonstrates how inference with ESMFold can be enhanced with ProteinTTT by adding just a few lines of code to enable customization. Next, Code snippet˜2 shows how ProteinTTT can be easily implemented for a PLM of interest by inheriting from the abstract TTTModule class. To integrate ProteinTTT within a model (e.g., ESM2), the user needs to implement methods that define the model’s vocabulary, an interface for predicting logits, and a specification of which modules need to be fine-tuned or remain frozen. The rest, i.e., the test-time training logic itself, is implemented within the unified TTTModule class.
Optimization.
We minimize the loss defined in Equation˜2 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (ruder2016overview). While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model (hayes2024simulating; lin2023evolutionary). Moreover, many models are pre-trained using the Adam optimizer (kingma2015adam) or its variants (loshchilov2019decoupled). However, it was shown that Adam results in less predictable behavior of test-time training compared to the SGD optimizer, possibly due to its more exploratory behavior (gandelsman2022test).
Customizing large models.
We aim for customization to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead. Since state-of-the-art models for many protein-oriented tasks are typically large, with up to billions of parameters, our aim presents two key challenges. First, when using pre-trained Transformers on a single GPU, even for the forward pass, the batch size is typically limited to only several samples due to the quadratic complexity of the inference (vaswani2017attention). Second, for the backward pass, even a batch size of one is not always feasible for large models. To address the first challenge, we perform forward and backward passes through a small number of training examples and accumulate gradients to simulate updates with any batch size. We address the second challenge by employing low-rank adaptation (LoRA; hu2021lora), which in practice enables fine-tuning of any model for which a forward pass on a single sample is feasible, due to a low number of trainable parameters. Section˜H.3 details how ESMFold (lin2023evolutionary), with its 3B-parameter ESM2 backbone $f$, can be efficiently customized, retaining its speed advantage while enhancing performance.
Appendix F Experimental details
In this section, we describe the proposed benchmark suite for the three customization tasks considered in this work: protein structure prediction (Section˜F.1), protein fitness prediction (Section˜F.2), and protein function prediction (Section˜F.3). Each subsection describes the application of ProteinTTT to the respective models, along with details on the data, metrics, and models. Table˜A3 additionally summarizes the hyperparameters used for the application of ProteinTTT to individual models.
F.1 Protein structure prediction
F.1.1 Datasets
@@
To evaluate the capabilities of ProteinTTT on protein structure prediction, we employ the CAMEO validation and test sets as described in lin2023evolutionary. Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^4] https://www.cameo3d.org/modeling (robin2021continuous) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022. Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (lin2023evolutionary). Therefore, we subselect the challenging validation and test sets where customization with ProteinTTT is suitable.
Specifically, we apply two standard criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (jumper2021highly), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors. Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in lin2023evolutionary). After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively.
F.1.2 Metrics
To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (zhang2004scoring) and LDDT (mariani2013lddt).
@@ F.1.3 Models
Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (hu2021lora) to all matrices involved in self-attention. This enables fine-tuning ESMFold + ProteinTTT on a single GPU.
@@
Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism for sampling multiple structure predictions. Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in lin2023evolutionary). To sample more predictions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids before passing them to the language model. Selecting the solution with the highest pLDDT may lead to improved predicted structure. Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + ProteinTTT, we employ the former as a baseline, running the method for the same number of steps.
HelixFold-Single.
HelixFold-Single is an MSA-free protein structure prediction model that combines representations from a pretrained protein language model with adapted AlphaFold2 geometric modules (EvoformerS and Structure) to directly predict atomic coordinates (fang2023method). We use the official implementation [^6] https://github.com/PaddlePaddle/PaddleHelix/tree/dev/apps/protein_folding/helixfold-single
HelixFold-Single + ProteinTTT.
DPLM2 Bit-based.
The DPLM2 Bit-based discrete diffusion protein language model (hsieh2025elucidating) extends DPLM2 by using bit-wise discrete modeling to enhance structure generation capabilities (wang2024dplm). DPLM2 is a multi-modal model that jointly models protein sequences and discretized structural tokens within a single discrete diffusion framework. In this work, we evaluate DPLM2 Bit-based on the task of structure prediction. Structure prediction is performed by initializing the structural tokens with masks and gradually denoising them based on the sequential tokens. We use the official implementation [^7] https://github.com/bytedance/dplm with the standard 650M-parameter model, 100 denoising steps, and the denoising strategy set to annealing@1.1:0.1.
DPLM2 Bit-based + ProteinTTT.
To apply ProteinTTT to DPLM2 Bit-based, we use the standard masked language modeling objective (Equation˜2). See Appendix˜B for further discussion. Please also note that we do not use confidence function $c$ with DPLM2 Bit-based as it does not implement pLDDT or any other confidence function for protein structure prediction.
@@
Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (devlin2018bert), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens). The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (razavi2019generating) operating on the protein geometry. In our experiments, we use the smallest, publicly available version of the ESM3 model (ESM3_sm_open_v0) [^8] https://github.com/evolutionaryscale/esm.
@@
F.2 Protein fitness prediction
F.2.1 Datasets
@@
ProteinGym [^9] https://github.com/OATML-Markslab/ProteinGym is the standard benchmark for protein fitness prediction (notin2024proteingym). The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins. We focus on the well-established zero-shot setup of the benchmark and do not experiment with the supervised setup, as it has not yet been fully incorporated into the official codebase at the time of this study. In total, the dataset contains 2.5M mutants with annotated ground-truth fitness. Since ProteinGym does not contain a data split for the zero-shot setup, employed in this work, we use the whole dataset as the test set.
@@
Figure A3: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work. A) MaveDB, mined from esposito2019mavedb, includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset. This is largely because most MaveDB assays post-filtering date to 2024, whereas the latest assays in ProteinGym date to 2023. B, C, D) MaveDB is of sufficient quality for model evaluation. Representative baselines, ESM2 and SaProt with both 35 million and 650 million parameters, evaluated on ProteinGym generalize effectively to MaveDB, following a similar distribution of predictions. Panel D illustrates the random subset of 50 proteins used for hyperparameter tuning for fitness prediction. Each point in the plots represents one protein and shows the Spearman correlation averaged across all assays corresponding to the protein (typically one assay per protein). The box plots standardly depict quartiles, medians, and outliers.
To establish a validation set disjoint from ProteinGym (notin2024proteingym), we mined MaveDB [^10] https://www.mavedb.org (esposito2019mavedb). As of August 1, 2024, the database contains 1178 Multiplexed Assays of Variant Effects (MAVEs), where each assay corresponds to a single protein, measuring the experimental fitness of its variants. We applied quality control filters to remove potentially noisy data. Specifically, we ensured that the UniProt identifier (uniprot2023uniprot) is valid and has a predicted structure available in the AlphaFold DB (varadi2022alphafold). We also excluded assays with fewer than 100 variants, as well as those where at least one mutation had a wrongly annotated wild type or where most mutations failed during parsing. Additionally, to ensure no overlap between datasets, we removed any assays whose UniProt identifier matched with those in ProteinGym, ensuring that the validation and test sets contain different proteins.
@@ F.2.2 Metrics
When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (notin2024proteingym). First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels. Then, we average the scores across five categories of assayed phenotypes, measuring the effects of mutations: catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
@@
Following (notin2024proteingym), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table˜A4). Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from hopf2017mutation: “Low” is defined as $N_{eff}/L<1$, “Medium” as $1<N_{eff}/L<100$, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
F.2.3 Models
@@
The ESM2 model is a bidirectional, BERT-like (devlin2018bert) Transformer trained on millions of protein sequences using masked modeling (lin2023evolutionary). The goal of protein fitness prediction is to predict the effects of mutations, and PLMs are often adapted to this task using zero-shot transfer via log odds ratio (notin2024proteingym; meier2021language). Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
@@
where the sum iterates over mutated positions $i\in T$ with $p(x^{m}_{i}|x_{\setminus i})$ and $p(x_{i}|x_{\setminus i})$ denoting the predicted probabilities of the mutated amino acid and the original one (i.e., wild type), respectively. The conditionals $x_{\setminus i}$ indicate that the input sequence to the model has the position $i$ masked. In this setup, the native (unmutated) sequence, where $T=\emptyset$, has a predicted fitness of 0. Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations. We follow the ProteinGym benchmark and use this formula (notin2024proteingym) to evaluate the fitness prediction capabilities of ESM2. We use the implementation of ESM2 from ProteinGym.
@@
We also experiment with a structure-aware protein language model, SaProt (su2023saprot). SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures. Specifically, SaProt uses the same Transformer architecture but expands its vocabulary by combining the 20 standard amino acid tokens with 20 structural tokens from the 3Di vocabulary, increasing the total alphabet size to 400. The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (razavi2019generating), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van2022foldseek).
Since SaProt is also a protein language model, it also uses Equation˜6 to score variants. However, please note that SaProt, as implemented in ProteinGym (notin2024proteingym), uses a slightly different version of the log odds ratio. In SaProt, the conditions in the log probabilities in Equation˜6 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions. During customization with ProteinTTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup. We use the implementation of SaProt from ProteinGym9.
@@
We experiment with the state-of-the-art fitness predictor, ProSST (li2024prosst). ProSST primarily improves upon SaProt (su2023saprot) by incorporating a larger vocabulary of structural tokens and employing disentangled attention mechanisms. Instead of relying on the 3Di alphabet optimized for protein structure search with Foldseek (van2022foldseek), li2024prosst pre-train a new autoencoder to denoise corrupted protein backbones and cluster the resulting latent space using the $K$-means algorithm (lloyd1982least). Notably, optimal performance for fitness prediction is achieved with $K=2048$ tokens, compared to just 20 in the 3Di vocabulary used by SaProt. We adopt this model in our experiments. Additionally, disentangled attention in ProSST enhances information propagation between sequence and structure within its Transformer blocks, further improving prediction performance. The model has 110M parameters in total.
ProSST, similarly to ESM2 and SaProt, is pre-trained using masked language modeling applied to protein sequence tokens. To score mutations on the ProteinGym benchmark (notin2024proteingym), ProSST also uses the log-odds ratio, but in a slightly different way compared to ESM2 and SaProt. Specifically, ProSST performs a single forward pass to predict log probabilities, which are then used to score all mutations. Formally, this approach modifies the log probability condition in Equation˜6, replacing $x_{\setminus i}$ with $x$.
@@
ProGen2.
For fitness prediction, we additionally experiment with one of the major autoregressive protein language models, ProGen2 (nijkamp2023progen2). Specifically, we experiment with ProGen of two sizes: ProGen2-small (151M parameters) and ProGen2-large (2.7B parameters). We obtain the pre-trained weights from the official GitHub repository [^11] https://github.com/salesforce/progen/tree/main/progen2. For ProGen2-large inference, we use floating-point 16 precision for computational efficiency.
ProGen2 + ProteinTTT.
To demonstrate the applicability of ProteinTTT in an autoregressive setting, we apply it to the ProGen2 (nijkamp2023progen2) language model. To perform the customization, we use the standard next-token prediction objective on a single target protein, following Equation˜4. Please see Appendix˜B for details.
@@
Finally, we experiment with MSA Transformer for fitness prediction (rao2021msa). Similar to ESM2 (lin2023evolutionary), MSA Transformer is pre-trained on large protein sequence datasets; however, it is trained on multiple sequence alignments (MSAs) rather than individual sequences.
Since MSA Transformer is also a protein language model, it can be used for fitness prediction in the same way as ESM2, as discussed above, by computing the log-odds ratio over the first sequence in the MSA in this case. We reproduce the results of MSA Transformer on the ProteinGym benchmark with two modifications: (1) we sample a weighted subset of 32 sequences from each MSA instead of 400, and (2) we use only one random seed instead of five for ensembling. These changes significantly reduce computational time while also slightly improving performance compared to the results reported in ProteinGym. This improvement may be explained by the fact that the performance of MSA Transformer saturates with increasing MSA input size (Figure 4 in rao2021msa).
@@
F.3 Protein function prediction
F.3.1 Datasets